OpenStudy (anonymous):

please someone explain to me on how to find the derivative of y= 4sec^2x

OpenStudy (ash2326):

@ICECREAM_HERO do you know basic differentiation?

OpenStudy (anonymous):

yup I just don't know how my teacher used the chain rule on this to get the answer o.O

OpenStudy (ash2326):

ok, it'd work like this \[\frac{d}{dx} (f(g(x))=f'(x)\times g'(x)\] I'll illustrate with an example \[\frac{d}{dx} (\sin ^2x)=2\times \sin x\times \frac{d}{dx} ( \sin x)\] first I found derivative of the square term, works the same as x^2=> 2x then I multiplied it with the derivative of sin x did you understand this?

OpenStudy (anonymous):

yup

OpenStudy (ash2326):

could you apply this method to your question?

OpenStudy (anonymous):

I'll try, I know how to use the chain rule in other problems but this one was just weird. The end answer is suppose to be \[8\sec^2xtanx\] but I keep getting something different

OpenStudy (ash2326):

You'd get this only \[\frac{d}{dx}4\times \sec^2 x\] \[4\frac{d}{dx} \sec^2 x\] Let sec x= f(x) then we have \[4\frac d {dx} (f(x))^2\] \[4 \times (2 f(x))\frac d {dx} f(x)\] now put f(x) 's value \[4\times (2\times \sec x)\frac {d}{dx} (\sec x)\] do you get this?

OpenStudy (anonymous):

sort of so on the last equation at the last part is that just taking the derivative of sec x? and how did you get \[(d/dx) 4 * \sec^2x\] sorry this is why I need help

OpenStudy (ash2326):

I chose f(x)= sec x so we needed to find \[\frac {d}{dx } (f(x))^2\] The first step is just like finding derivative of x^2=> 2x it's multiplied by the derivative of f(x) i.e f'(x) \[2 f(x) f'(x)\]

OpenStudy (anonymous):

OHHHH otay I see o.o

OpenStudy (ash2326):

Do you get it through?

OpenStudy (anonymous):

better than before it's just a different teaching style so how would you get the tanx in the solution

OpenStudy (ash2326):

derivative of sec x is \[\frac{d}{dx} (\sec x)=\sec x\times \tan x\]

OpenStudy (anonymous):

HAHA OHH never mind sorry that was on my formula sheet XD should've known that and thank you so much!! I actually understand this better

OpenStudy (ash2326):

Welcome :D